# Approximating Feynman Integrals Using Complete Monotonicity and Stieltjes Properties

### v1

@@ 3.6 Numerical evaluation of a 20-loop ‘banana’ integral
The equal mass $L$-loop ‘banana’ integral $I_{L}(x)$ in $D=2$ admits a representation as a one-fold Bessel integral in the Euclidean region, i.e. for $x\geq-(L+1)^{2}$, see refs. Groote:2005ay; Pogel:2022vat,
@@
where $J_{0}(z)$ and $K_{0}(z)$ are the Bessel function of the first kind and the modified Bessel function of the second kind, respectively.
@@
We have omitted the overall factor of $2^{L}$ from our analysis to keep the numbers obtained manageable. These Bessel moments have been computed in closed form for $L=1,2,3$, see ref. Bailey:2008ib. To achieve a reasonable precision for the $20$-loop ‘banana’ integral, we compute numerically the first $20$ Bessel moments in eq. (36), to an accuracy of $50$ digits i.e., $n=20$ and $k=50$. This takes roughly $35$ s. We then use this information to obtain the corresponding Padé approximants.
@@ 4 Discussion and outlook
A related approach, followed in ref. Zeng:2023jek, is to reconstruct the terms in the Laurent series based on numerically sampling the results for specific values of $\epsilon$. The Stieltjes property opens up another possibility, which is to first obtain a Padé approximation, for general $\epsilon$. We have found promising evidence that the latter can be expanded in $\epsilon$ to obtain approximations of the coefficients in the Laurent expansion. We leave this fascinating direction for future work.
Although this work focused on Feynman integrals, we expect our method to have broader applications, including in cosmology and gravitational physics. This expectation comes from the fact that, as shown in ref. Henn:2024qwe, several objects appearing in these areas–such as cosmological correlators and Euler–Mellin integrals–satisfy the CM property.
@@
  • Numerical Padé interpolation. There are various situations relevant to collider physics where Feynman integrals are not (yet) accessible analytically, but where numerical integration methods can produce results at certain phase-space points. This typically comes with high computational cost and long runtimes. In such a situation, the Stieltjes property justifies using a multi-point Padé interpolation. Our method could also be combined with other lattice reconstruction methods, such as those applied recently in ref. Barrera:2025uin.

### v2

@@ 3.6 Numerical evaluation of a 20-loop ‘banana’ integral
The equal mass $L$-loop ‘banana’ integral $I_{L}(x)$ in $D=2$ admits a representation as a one-fold Bessel integral in the Euclidean region, i.e. for $x\geq-(L+1)^{2}$, see refs. Groote et al. (2007); Pögel et al. (2023),
@@
where $J_{0}(z)$ and $K_{0}(z)$ are the Bessel function of the first kind and the modified Bessel function of the second kind, respectively. The Bessel integral representation can be used to derive the Picard–Fuchs differential equations satisfied by these integrals to very high loop orders; this has been carried out up to $L=20$, see refs. Vanhove (2014); de la Cruz and Vanhove (2024).
@@
We have omitted the overall factor of $2^{L}$ from our analysis to keep the numbers obtained manageable. These Bessel moments have been computed in closed form for $L=1,2,3$, see ref. Bailey et al. (2008). To achieve a reasonable precision for the $20$-loop ‘banana’ integral, we compute numerically the first $20$ Bessel moments in eq. (36), to an accuracy of $50$ digits i.e., $n=20$ and $k=50$. This takes roughly $35$ s. We then use this information to obtain the corresponding Padé approximants.
@@ 4 Discussion and outlook
A related approach, followed in ref. Zeng (2023), is to reconstruct the terms in the Laurent series based on numerically sampling the results for specific values of $\epsilon$. The Stieltjes property opens up another possibility, which is to first obtain a Padé approximation, for general $\epsilon$. We have found promising evidence that the latter can be expanded in $\epsilon$ to obtain approximations of the coefficients in the Laurent expansion. We leave this fascinating direction for future work.
Although this work focused on Feynman integrals, we expect our method to have broader applications, including in cosmology and gravitational physics. This expectation comes from the fact that, as shown in ref. Henn and Raman (2025), several objects appearing in these areas–such as cosmological correlators and Euler–Mellin integrals–satisfy the CM property.
@@
  • Numerical Padé interpolation. There are various situations relevant to collider physics where Feynman integrals are not (yet) accessible analytically, but where numerical integration methods can produce results at certain phase-space points. This typically comes with high computational cost and long runtimes. In such a situation, the Stieltjes property justifies using a multi-point Padé interpolation. Our method could also be combined with other lattice reconstruction methods, such as those applied recently in ref. Barrera et al. (2025).